# Quadratic dispersion relations in gapless frustration-free systems

### v1

@@ VI Conclusion
In this paper, we proposed two conjectures regarding frustration-free systems. The first conjecture posited that no finite-size gap opens between the degenerate ground states of frustration-free systems. While this has been implicitly assumed in previous studies Bravyi and Gosset (2015); Gosset and Mozgunov (2016); Gosset and Huang (2016), we pointed out the necessity of being cautious as it is actually a non-trivial assumption.
The second conjecture stated that in a gapless frustration-free system with translational symmetry, the dispersion relation of low-energy excitations near a certain wave number $\bm{k}_{0}$ can be bounded above by a quadratic dispersion, i.e., $O\big{(}|\bm{k}-\bm{k}_{0}|^{2},L^{-2}\big{)}$. We proved this conjecture for the case of an $s=1/2$ spin chain with nearest-neighbor interactions by combining the result of a seminal work by Bravyi and Gosset Bravyi and Gosset (2015) and the min-max theorem. Furthermore, we extended these results to two- and three-dimensional models on the cubic lattice. However, it became clear that this method could not be directly extended to cases where three or more neighboring spins interact simultaneously. Extending our results to cases where the Hilbert space dimension at each site is larger than two or to models with longer-range interactions constitutes important future work.
Finally, there is a related, more general conjecture that the finite-size gap $\epsilon=E_{\tilde{N}_{\mathrm{deg}}+1}$ for a gapless frustration-free Hamiltonian can be bounded above by $O(L^{-2})$ regardless of the presence or absence of translational symmetry or the details of the boundary conditions. In fact, in the accompanying paper Masaoka et al. (2024), we prove this statement for critical frustration-free systems in which a ground-state correlation function shows a power-law behavior. However, such an argument is not applicable, for example, to the uncle Hamiltonian for the GHZ state since all correlation functions decays exponentially as discussed in Sec. V.2.
@@ Acknowledgements.
We thank Hosho Katsura for useful discussions. The work of H.W. is supported by JSPS KAKENHI Grant No. JP24K00541. This research is funded in part by the Gordon and Betty Moore Foundation’s EPiQS Initiative, Grant GBMF8683 to T.S.
Appendix A Proof of the inequality (11)
@@
$\max_{|\Phi\rangle\in\mathcal{M}_{j}}\langle\Phi|\hat{H}^{\prime}|\Phi\rangle$ $=\langle\Phi_{*}^{\prime}|\hat{H}^{\prime}|\Phi_{*}^{\prime}\rangle$
$\geq\langle\Phi_{*}|\hat{H}^{\prime}|\Phi_{*}\rangle$
@@
This relation holds for any $\mathcal{M}_{j}\subset\mathcal{H}_{D}$. Hence, we arrive at the inequality (11).
@@
$\hat{U}|1\rangle=e^{i\theta}(-u_{2}^{*}|0\rangle+u_{1}^{*}|1\rangle),$ (95)
@@
Hence
@@
$=e^{-2i\theta}\begin{pmatrix}\langle\psi|(u_{1}^{*}|0\rangle-u_{2}|1\rangle)(u_{2}^{*}|0\rangle+u_{1}|1\rangle)&\langle\psi|(u_{2}^{*}|0\rangle+u_{1}|1\rangle)(u_{2}^{*}|0\rangle+u_{1}|1\rangle)\\ -\langle\psi|(u_{1}^{*}|0\rangle-u_{2}|1\rangle)(u_{1}^{*}|0\rangle-u_{2}|1\rangle)&-\langle\psi|(u_{2}^{*}|0\rangle+u_{1}|1\rangle)(u_{1}^{*}|0\rangle-u_{2}|1\rangle)\end{pmatrix}$
$=e^{-2i\theta}\begin{pmatrix}u_{1}&-u_{2}^{*}\\ u_{2}&u_{1}^{*}\end{pmatrix}\begin{pmatrix}\langle\psi|(u_{1}^{*}|0\rangle-u_{2}|1\rangle)|1\rangle&\langle\psi|(u_{2}^{*}|0\rangle+u_{1}|1\rangle)|1\rangle\\ -\langle\psi|(u_{1}^{*}|0\rangle-u_{2}|1\rangle)|0\rangle&-\langle\psi|(u_{2}^{*}|0\rangle+u_{1}|1\rangle)|0\rangle\end{pmatrix}$
@@
Therefore, by choosing $U=\begin{pmatrix}u_{1}&-u_{2}^{*}\\ u_{2}&u_{1}^{*}\end{pmatrix}$ and $\theta$ properly, one gets the form in Eq. (25), which is equivalent to the parametrization in Eq. (20).
Appendix C The case of $\det m_{\psi}=0$.
@@
where $|v^{\perp}\rangle=i\gamma|0\rangle+\delta|1\rangle$ with $\gamma^{2}+\delta^{2}=1$. According to Ref. Bravyi and Gosset (2015), excitations under OBC are gapped in this case, which implies that excitations under PBC are also gapped.
@@
where $|v\rangle=\delta|0\rangle+i\gamma|1\rangle$ is orthogonal to $|v^{\perp}\rangle$. Translation-invariant product states $|v\cdots v\rangle$ and $|0\cdots 0\rangle$ are also ground states but only the former is orthogonal to the domain-wall states and the latter is not linearly independent. In contrast, under PBC, domain wall states violate the boundary term and the two translation-invariant product states give the ground states. Thus the formula in Eq. (32) is applicable to this case.

### v2

@@ VI Conclusion
In this work, we proposed two conjectures regarding FF systems. The first one posited that no finite-size splitting opens between the degenerate ground states. While this has been implicitly assumed in previous studies Bravyi and Gosset (2015); Gosset and Mozgunov (2016); Gosset and Huang (2016), we pointed out the necessity of being cautious as it is actually a non-trivial assumption. The second conjecture stated that in a gapless FF system with translational symmetry, the dispersion relation of low-energy excitations near a certain wave number $\bm{k}_{0}$ can be bounded above by a quadratic dispersion. We proved this for the case of $s=1/2$ spin models with nearest-neighbor interactions defined on the hyper cubic lattice. However, this method could not be directly extended to models with longer-range interactions or larger local Hilbert space dimensions. Extending our results to these cases constitutes important future work.
Finally, there is a related, more general conjecture that the finite-size gap $\epsilon\coloneqq E_{\tilde{N}_{\mathrm{deg}}+1}$ for gapless FF Hamiltonians can be bounded above by $O(L^{-2})$ regardless of the presence or absence of translational symmetry or the details of the boundary conditions. In fact, in the accompanying paper Masaoka et al. (2024), we prove this statement for critical FF systems in which a ground-state correlation function shows a power-law behavior. However, such an argument is not applicable to the uncle Hamiltonian for the GHZ state since all correlation functions decays exponentially.
@@ Acknowledgements.
We thank Hosho Katsura and Zijian Wang for useful discussions. The work of H.W. is supported by JSPS KAKENHI Grant No. JP24K00541. This research is funded in part by the Gordon and Betty Moore Foundation’s EPiQS Initiative, Grant GBMF8683 to T.S.
Appendix A Proof of the min-max principle
@@
$\max_{|\Phi\rangle\in\mathcal{M}_{j}}\langle\Phi|\hat{H}^{\prime}|\Phi\rangle$ $=\langle\Phi_{*}^{\prime}|\hat{H}^{\prime}|\Phi_{*}^{\prime}\rangle\geq\langle\Phi_{*}|\hat{H}^{\prime}|\Phi_{*}\rangle$
@@
This relation holds for any $\mathcal{M}_{j}\subset\mathcal{H}_{D}$. Hence, we arrive at the theorem.
@@
$\hat{U}|1\rangle=e^{i\theta}(-u_{2}^{*}|0\rangle+u_{1}^{*}|1\rangle).$ (46)
@@
Under this local unitary transformation, the matrix $m_{\psi}$ defined by
$m_{\psi}\coloneqq\begin{pmatrix}\langle\psi|01\rangle&\langle\psi|11\rangle\\ -\langle\psi|00\rangle&-\langle\psi|10\rangle\end{pmatrix}$ (49)
is changed to
@@
Therefore, by choosing $U=\begin{pmatrix}u_{1}&-u_{2}^{*}\\ u_{2}&u_{1}^{*}\end{pmatrix}$ and $\theta$ properly, one gets the form
$m_{\psi}=\begin{pmatrix}\alpha-i\beta&\delta\\ 0&-\alpha+i\gamma\end{pmatrix}.$ (51)
which is equivalent to the parametrization in Eq. (6) of the main text.
Appendix C Derivation of Eq. (16)
@@
where $|v^{\perp}\rangle=i\gamma|0\rangle+\delta|1\rangle$ with $\gamma^{2}+\delta^{2}=1$. According to Ref. Bravyi and Gosset, 2015, excitations under OBC are gapped in this case, which implies that excitations under PBC are also gapped.
@@
where $|v\rangle=\delta|0\rangle+i\gamma|1\rangle$ is orthogonal to $|v^{\perp}\rangle$. Translation-invariant product states $|v\cdots v\rangle$ and $|0\cdots 0\rangle$ are also ground states but only the former is orthogonal to the domain-wall states and the latter is not linearly independent. In contrast, under PBC, domain wall states violate the boundary term and the two translation-invariant product states give the ground states.